Automobile air bags are inflated by nitrogen gas generated by the rapid decomposition of sodium azide. If an air bag has a volume of 36 L and is to be filled with nitrogen gas at 1.15 atm and 26.0OC, how many grams of NaN3 must be decomposed?


NaN3(s) = Na(s) + N2(g)


1
Expert's answer
2021-11-05T01:07:00-0400

Using PV=nRT

n=PV/RT

n=(1.15×36)/0.082×299

n=1.69moles

Mole=mass/molar mass

Mass=mole × molar mass

Mass=1.69×65

Mass=109.85g